# OPENXRD: A Comprehensive Benchmark and Enhancement Framework for LLM/MLLM XRD Question Answering

### v1

@@ 2.4 Input Context Structuring
We implement a structured input preparation approach via a fusion module. For open-book evaluations, we employ a retrieve-then-read prompt structure where the supporting material precedes the question:
@@
where $x_{s}$ is the supporting textual content, $x_{q}$ represents the question text, and $x_{o}$ indicates the provided answer choices. The Format function organizes these components with clear delimiters and instructional guidance. For closed-book evaluations, the supporting material is omitted ($x_{s}=\emptyset$), while the remaining formatting is preserved.
This structured approach aligns with established retrieval-augmented generation paradigms, wherein external knowledge serves as contextual reference material. Figure 4 demonstrates the architecture of this open-book QA pipeline. Currently, our implementation is exclusively textual; however, this framework can be extended to multimodal data (such as XRD patterns and crystallographic diagrams) as generative vision models become sufficiently advanced.
@@ 2.5 Evaluation Metrics
Accuracy=|{correctly answered questions}||{all $\text{Accuracy}=\frac{|\{\text{correctly answered questions}\}|}{|\{\text{all questions}\}|}\times 100\%$ (2)
@@
where $Q_{t}$ is the set of questions associated with subtask $t$, $y_{q}$ is the correct answer, and $\bm{M}(q)$ denotes the model’s prediction.
To assess the efficacy of open-book augmentation, we measure performance improvement $\Delta$ as:
@@ 3 Results
In this section, we present a detailed evaluation of our open-book QA system for crystallography, using a curated set of 217 XRD-related questions. Each question can be answered in two ways: Closed-book mode, where the model relies solely on its internal knowledge, and Open-book mode, where it consults supporting textual material. Our goal is to determine to what extent supporting textual material access bolsters model performance, particularly for advanced domain-specific queries. We conduct our experiments in two phases: first using AI-generated supporting materials provided by GPT-4.5, and then using expert-reviewed versions of these materials to assess whether human domain expertise further enhances performance.
@@ 3.1 Setup and Baselines
Models Compared. We evaluate several large language (and vision-language) models. This includes the O-family Models such as O3-mini (O3-mini-2025-01-31) and O1 (O1-2024-12-17), which are designed for scientific reasoning tasks. We also look at various GPT-4 Variants like GPT-4.5-preview (GPT-4.5-preview-2025-02-27), GPT-4-turbo-2024-04-09, GPT-4-preview-0125, and GPT-4-0613. These versions were tested in an interim stage that incorporates "supporting textual material" into GPT-centered language processing. Additionally, we explore pipelines utilizing LLaVA, including LLaVA-v1.6-mistral-7B, LLaVA-v1.5 (7B, 13B), and LLaVA-onevision-QWEN2, which combine a vision encoder (ViT) with an LLM backbone. These models are significant as they use text-only configurations with "supporting textual materials" that include image captions but not the images themselves, within the broader context of this research.
@@
Inference Setup. We employ a zero-shot inference paradigm for all models without parameter updating or fine-tuning on our crystallography evaluation set. For each query, we construct a prompt consisting of the question and enumerated options, formatted consistently across all models. No in-context learning examples or demonstrations are provided, requiring models to rely entirely on their parametric knowledge. All foundation models are evaluated using their published weights and architectures; O-family models have documented parameter-efficient training on scientific corpora, while vision-language models in the LLaVA family were primarily trained on general domain text-image pairs without domain-specific adaptation to crystallography.
@@
Experimental Progression. Our experimental approach follows two distinct phases: In the first phase, we use GPT-4.5 to generate supporting textual materials for each question, carefully prompting it to provide relevant information without revealing the answer. Following this, in the second phase, we engaged three Ph.D. students specializing in crystallography, with 4-7 years of research experience, to review and refine the supporting materials for improved accuracy and pedagogical clarity. This two-phase approach allows us to quantify the value added by human domain expertise in the construction of supporting materials relative to state-of-the-art AI generation.
@@ 3.2 Closed-Book Mode Observations
Table 1 presents the closed-book mode accuracy achieved on our 217-question crystallography benchmark, and Figure 5 illustrates the categorized closed-book performance of the models. O3-mini performs extremely well, achieving about 89% accuracy, even surpassing some GPT-4 variants, likely due to targeted domain pre-training. However, smaller LLaVA-based models face challenges with advanced reflection extinctions, often incorrectly identifying which plane indices vanish for BCC or confusing zone-axis notation. Despite these issues, nearly all models correctly answer basic factual questions like the number of crystal systems and who discovered X-ray diffraction.
@@
Table 1: Comprehensive "closed-book mode" accuracy results on the 217-item crystallography QA benchmark, including older GPT-4 variants and newly reported models.
Rank Model (Closed-book mode) Accuracy (%) Correct #
1 GPT-4.5-preview 93.09 202/217
2 O3-mini 88.94 193/217
3 O1 87.56 190/217
4 GPT-4-turbo-preview 83.41 181/217
5 GPT-4 81.11 176/217
6 LLaVA-v1.6-34B 66.80 145/217
7 LLaVA-onevision-QWEN2-7B-si 66.40 144/217
8 LLaVA-onevision-QWEN2-7B-ov-chat 65.90 143/217
9 LLaVA-onevision-QWEN2-7B-ov 65.40 142/217
10 LLaVA-v1.6-mistral-7B 53.00 115/217
11 LLaVA-onevision-QWEN2-0.5B-ov 47.50 103/217
12 LLaVA-v1.5-13B 46.50 101/217
13 LLaVA-onevision-QWEN2-0.5B-si 46.10 100/217
14 LLaVA-v1.5-7B 18.00 39/217
Large ModelsMid-sized ModelsSmall Models88.9%66.2%37.2%Model GroupsClosed-book performance
@@
Fig.  5: Comparison of closed-book performance by model size group. Large Models include GPT-4 variants and O3-mini; Mid-sized Models include LLaVA-34B and QWEN2-7B variants; Small Models include LLaVA-13B, LLaVA-7B, and QWEN2-0.5B variants.

### v2

@@ 2.4 Input Context Structuring
We implement a structured input preparation approach via a fusion module. For open-book evaluations, we employ a retrieve-then-read style input format (without executing retrieval) where the supporting material precedes the question:
@@
where $x_{s}$ is the supporting textual content, $x_{q}$ represents the question text, and $x_{o}$ indicates the provided answer choices. The Format function organizes these components with clear delimiters and instructional guidance. For closed-book evaluations, the supporting material is omitted ($x_{s}=\emptyset$), while the remaining formatting is preserved. This structured approach aligns with established retrieval-augmented generation paradigms, wherein external knowledge serves as contextual reference material. Here, $x_{s}$ is an oracle helper passage selected and refined offline; no retriever, corpus chunking, or ranking is used in OPENXRD. Figure 5 demonstrates the architecture of this open-book QA pipeline. Currently, our implementation is exclusively textual; however, this framework can be extended to multimodal data (such as XRD patterns and crystallographic diagrams) as generative vision models become sufficiently advanced.
@@ 2.5 Evaluation Metrics
$\text{Accuracy}=\frac{|\{\text{correctly answered questions}\}|}{|\{\text{all questions}\}|}\times 100\%$ (2)
@@
where $Q_{t}$ is the set of questions associated with subtask $t$, $y_{q}$ is the correct answer, and $\bm{M}(q)$ denotes the model’s prediction. To assess the efficacy of open-book augmentation, we measure performance improvement $\Delta$ as:
@@
2.6 Benchmark Reproducibility and Extension Guidelines
To establish OPENXRD as a reliable community benchmark with reproducible content standards, we provide documentation and tools for three primary use cases.
Public release. All benchmark materials are publicly available at https://github.com/niaz60/OpenXRD and will be archived on Zenodo with DOI upon acceptance. The release includes: full question set with answers and explanations; AI-generated and expert-reviewed supporting passages for all 217 questions; GPT-4.5 generation prompts with parameters; expert-review guidelines with the four-dimension quality rubric; token matching scripts; and validation protocols for answer-guidance testing.
Evaluating new models. Researchers should download the full benchmark, run the closed-book and open-book protocols (AI and expert passages) using our scripts, and compare against baseline tables to contextualize results.
Creating new crystallography questions. Follow our question template (prompt, 3–4 options, single correct answer, brief explanation, subtask label); generate initial passages with our GPT-4.5 prompt template (“Given question… provide a concise explanation… Target 500–600 tokens”); regenerate any passage that reveals the answer or has technical errors; apply expert review using the four dimensions; token-match AI and expert versions to <5% difference ($r>0.90$).
Domain adaptation. For new scientific domains, adapt the generation prompt to domain terminology, recruit domain experts to review using the same four dimensions, validate answer-guidance properties on held-out models, and document domain-specific modifications.
Validation protocol. To ensure supporting materials are answer-guiding without being answer-revealing: (1) choose 3–5 held-out models across capacity tiers; (2) test on 20–30 questions in three conditions—closed-book, explicit-answer materials, candidate materials; (3) candidate materials pass if performance lies between closed-book and explicit-answer conditions and preserves the closed-book model ranking. This confirms that materials aid reasoning without providing shortcuts.
2.7 Context Assimilation vs. Parameter Fine-Tuning
OPENXRD evaluates context assimilation—how models utilize short, answer-guiding passages provided at inference time. We intentionally do not modify model weights through parameter-efficient fine-tuning (PEFT) methods such as LoRA 39, as this would target a fundamentally different research question: parametric knowledge embedding rather than inference-time guidance utilization. Throughout this work, expert refinement refers to refinement of the supporting passages, not to any modification of evaluated LLMs. Accordingly, any performance changes in open-book mode should be interpreted as improved use of provided context (context assimilation), not weight-level refinement of model reasoning.
Several factors motivate this scope decision. First, many of the 74 models in our evaluation pool are API-only or closed-weight systems (GPT-4.5, O3-mini, Claude variants) that cannot be fine-tuned uniformly, making fair PEFT comparisons intractable. Second, our 217-question benchmark is intentionally reserved as a held-out evaluation set to preserve benchmark integrity; fine-tuning on these questions or near-duplicates would introduce data leakage and undermine reproducibility. Third, PEFT optimization conflates parametric adaptation effects with the inference-time assimilation dynamics we aim to isolate.
OPENXRD is complementary to PEFT research rather than competitive with it. Researchers can fine-tune open-weight models on external crystallographic corpora and subsequently use OPENXRD to test whether inference-time guidance remains beneficial, becomes redundant, or introduces interference after domain adaptation. This decomposition of parametric knowledge versus contextual reasoning is precisely the diagnostic capability OPENXRD provides. For researchers aiming to perform PEFT studies using our benchmark, we provide the complete dataset, evaluation scripts, and code framework on our project repository. This open-source release enables researchers to design their own PEFT experiments using standard techniques such as LoRA with appropriate data splits to prevent leakage between training and evaluation sets.
@@ 3 Results
In this section, we present a detailed evaluation of our OPENXRD benchmark for crystallography, using a curated set of 217 XRD-related questions. Each question can be answered in two ways: Closed-book mode, where the model relies solely on its internal knowledge, and Open-book mode, where it consults supporting textual material. Our goal is to determine to what extent supporting textual material access bolsters model performance, particularly for advanced domain-specific queries. We conduct our experiments in two phases: first using AI-generated supporting materials provided by GPT-4.5, and then using expert-reviewed versions of these materials to assess whether human domain expertise further enhances performance.
@@ 3.1 Setup and Baselines
3.1.1 Models Compared
We conduct a comprehensive evaluation across 74 state-of-the-art language models and vision-language models spanning multiple architectural families, parameter scales, and specialization domains. This extensive model selection enables systematic analysis of how model architecture, scale, and domain adaptation affect crystallographic reasoning performance under both closed-book and open-book conditions.
OpenAI models. We evaluate 13 OpenAI models across multiple generations and capabilities. The next-generation GPT-5 series includes gpt-5 and gpt-5-codex74, 75. The reasoning-optimized O-series comprises o3-mini (o3-mini-2025-01-31), o1 (o1-2024-12-17), and o1-mini76, 72. The GPT-4 family includes gpt-4.5-preview (gpt-4.5-preview-2025-02-27), gpt-4o, gpt-4o-mini, gpt-4-turbo (gpt-4-turbo-2024-04-09), gpt-4-turbo-preview (gpt-4-preview-0125), gpt-4-0314 (gpt-4-0613), and the base gpt-473, 71, 70, 68. We also evaluate earlier-generation models gpt-3.5-turbo and gpt-3.5-turbo-16k69, 67. Anthropic Claude models. We assess four Claude variants spanning two generations: claude-3.5-sonnet, claude-3.5-haiku, claude-3-opus, and claude-3-haiku9, 7, 8. Meta LLaMA models. The LLaMA family is represented by six models across three generations and multiple parameter scales: llama-3.1-405b-instruct (405B parameters), llama-3.1-70b-instruct and llama-3-70b-instruct (70B parameters), llama-3.1-8b-instruct and llama-3-8b-instruct (8B parameters), and llama-3.2-3b-instruct (3B parameters)54, 57.
QWEN/Alibaba models. We evaluate 10 models from Alibaba’s QWEN ecosystem. General-purpose models include qwen3-next-80b-a3b-instruct, qwen3-next-80b-a3b-thinking (reasoning-optimized variant), qwen-2.5-72b-instruct, qwen-plus, qwen-max, and qwen-2.5-7b-instruct105, 85. Code-specialized variants include qwen3-coder-flash, qwen3-coder-plus, and qwen-2.5-coder-32b-instruct94, 40. We also evaluate tongyi-deepresearch-30b-a3b, Alibaba’s research-focused model95. Mistral AI models. Seven Mistral variants are evaluated: mistral-large and mistral-small (dense models)61, 63, mistral-7b-instruct and mistral-7b-instruct-v0.1 (7B base models)42, the mixture-of-experts models mixtral-8x22b-instruct and mixtral-8x7b-instruct43, 62, and the multimodal pixtral-12b2. DeepSeek models. We evaluate two DeepSeek variants: deepseek-v3.1-terminus and deepseek-chat27. Google models. Three Google models are assessed: the flagship gemini-2.5-pro and the open-weight models gemma-2-27b-it and gemma-2-9b-it31, 32. Amazon Nova models. We evaluate Amazon’s Nova series across three capability tiers: nova-pro-v1, nova-lite-v1, and nova-micro-v15. X.AI models. We include X.AI’s grok-4-fast in our evaluation104. LLaVA vision-language models. Ten LLaVA variants are evaluated, spanning multiple generations and backbone architectures. These include llava-v1.6-34b and llava-v1.6-mistral-7b from the 1.6 series52, llava-v1.5-13b and llava-v1.5-7b from the 1.5 series51, and six LLaVA-OneVision models with QWEN2 backbones: llava-onevision-qwen2-7b-si, llava-onevision-qwen2-7b-ov-chat, llava-onevision-qwen2-7b-ov (7B variants), and llava-onevision-qwen2-0.5b-ov, llava-onevision-qwen2-0.5b-si (0.5B variants)49. These models combine vision encoders (ViT) with LLM backbones (Mistral, LLaMA, QWEN) for multimodal understanding. In our text-only evaluation setting, these models process supporting textual materials that may include image captions but not the actual images themselves. LLaMAT crystallography-specialized models: We evaluate three domain-adapted LLaMAT variants that have been pre-trained on crystallographic information files (CIF): llamat-3-chat, llamat-2-chat, and llamat-260. These models represent explicit attempts to enhance materials science reasoning through domain-specific pre-training.
Dziner models. We assess three Dziner-QWEN variants spanning different parameter scales: dziner-qwen-2.5-72b, dziner-qwen-2.5-coder-32b, and dziner-qwen-2.5-7b6.
Perplexity models. Two Perplexity variants are evaluated: sonar-pro and sonar81, 80. Additional specialized models. Our evaluation includes several community-developed and specialized models: Meituan’s longcat-flash-chat56, Arcee-AI’s afm-4.5b11, the community-tuned models undi95/remm-slerp-l2-13b and gryphe/mythomax-l2-13b98, 34, the vision-language models honeybee-13b and honeybee-7b17, and the routing system openrouter/auto77.
This comprehensive model selection, spanning from 0.5B to 405B+ parameters and covering generalist, reasoning-optimized, code-specialized, domain-adapted, and vision-language architectures, enables robust analysis of how different model capabilities interact with domain-specific supporting materials in crystallographic question answering.
Several models were considered but ultimately excluded from our evaluation due to architectural incompatibilities or technical limitations. The Ether-0 model65, designed specifically as a reward model for scoring the quality of scientific reasoning in chemistry rather than as a generative question-answering system, was deemed inappropriate for our benchmark. Reward models are trained using preference learning objectives to rank or score candidate responses, fundamentally differing from generative models that produce answers directly. This architectural mismatch makes Ether-0 incompatible with our multiple-choice QA evaluation framework, which requires models to generate or select answers rather than evaluate them.
We also encountered technical challenges with several LLaMAT variants—specialized adaptations of LLaMA models pre-trained on crystallographic information files (CIF). While three LLaMAT models (llamat-3-chat, llamat-2-chat, and llamat-2) were successfully evaluated after configuring max_model_len=2048 to accommodate their tokenization schemes, three additional variants (llamat-3-base, llamat-3-cif, and llamat-2-cif) exhibited incomplete HuggingFace configurations: the base llamat-3 model lacked proper architecture specifications in config.json, while both CIF-specialized variants had malformed model cards with missing tokenizer bindings that prevented inference initialization. Given that the successfully evaluated chat-optimized LLaMAT variants achieved moderate baseline performance (50.69-57.14% accuracy in closed-book mode; see Table 2), the exclusion of these broken variants does not materially impact our conclusions about crystallography-specialized models.
To facilitate systematic performance analysis across different capability levels, we categorize the 74 evaluated models into three size groups based on parameter count and architectural complexity, as summarized in Table 1. This categorization enables structured comparison of how model scale affects the ability to assimilate external domain knowledge in crystallography tasks.
Table 1: Model categorization by size for systematic performance analysis. The 74 evaluated models are grouped into three categories to examine scale-dependent effects on external knowledge assimilation in crystallography.
Category Size Range Representative Models
Small models << 10B parameters Mistral-7B, Phi-3-Mini (3.8B), Llama-3-8B,
LLaVA-v1.5-7B, QWEN2-0.5B, HoneyBee-7B,
LLaMAT-2 (7B), AFM-4.5B
Mid-sized models 10B–70B parameters Llama-3-70B, LLaVA-v1.6-34B, QWEN-2.5-32B,
[…]
Gemma-2-27B, Mistral-Small (24B)
Large models >> 70B parameters Llama-3.1-405B, QWEN-2.5-72B,
(or closed-weight dziner-qwen-2.5-72B, GPT-4, GPT-5,
API models) GPT-4.5, O3-mini, Claude-3-Opus,
Gemini-2.5-Pro, DeepSeek-V3.1
3.1.2 Inference Setup
@@
We employ a zero-shot inference paradigm for all models without parameter updating or fine-tuning on our crystallography evaluation set. For each query, we construct a prompt consisting of the question and enumerated options, formatted consistently across all models. No in-context learning examples or demonstrations are provided, requiring models to rely entirely on their parametric knowledge. All foundation models are evaluated using their published weights and architectures; O-family models have documented parameter-efficient training on scientific corpora, while vision-language models in the LLaVA family were primarily trained on general domain text-image pairs without domain-specific adaptation to crystallography.
3.1.3 Experimental Progression
@@
Our experimental approach follows two distinct phases: In the first phase, we use GPT-4.5 to generate supporting textual materials for each question, carefully prompting it to provide relevant information without revealing the answer. Following this, in the second phase, we engaged three Ph.D. students specializing in crystallography, with four to seven years of research experience, to review and refine the supporting materials for improved accuracy and pedagogical clarity. This two-phase approach allows us to quantify the value added by human domain expertise in the construction of supporting materials relative to state-of-the-art AI generation.
@@ 3.2 Closed-Book Mode Observations
Table 2 presents the closed-book mode accuracy achieved on our 217-question crystallography benchmark, and Figure 6 illustrates the categorized closed-book performance of the models.
@@
O3-mini performs extremely well, achieving about 93.55% accuracy, even surpassing some GPT-4 variants, possibly reflecting training or optimization details not disclosed; we report the observed performance. However, smaller LLaVA-based models face challenges with advanced reflection extinctions, often incorrectly identifying which plane indices vanish for BCC or confusing zone-axis notation. Despite these issues, nearly all models correctly answer basic factual questions like the number of crystal systems and who discovered X-ray diffraction.
@@
Table 2: Comprehensive "closed-book mode" accuracy results on the 217-item crystallography QA benchmark, including older GPT‑4 variants and newly reported models. The rightmost column shows model parameter sizes: B = billions, T = trillions, U = undisclosed. For Mixture-of-Experts (MoE) models, format is total/active parameters. Estimated values marked with $\sim$, N/A for routing services.
Rank Model (Closed-book mode) Acc. (%) Correct Params
1 openai/gpt-5 96.77 210/217 $\sim$300B
2 x-ai/grok-4-fast 96.31 209/217 U
2 openai/gpt-5-codex 96.31 209/217 $\sim$300B
4 google/gemini-2.5-pro 95.39 207/217 U
5 o3-mini 93.55 203/217 U
6 meituan/longcat-flash-chat 93.09 202/217 560B/27B
7 gpt-4.5-preview 92.63 201/217 U
8 anthropic/claude-3.5-sonnet 91.24 198/217 $\sim$175B
9 openai/gpt-4o 90.74 196/216 $\sim$200B
10 dziner-qwen-2.5-72b 90.32 196/217 $\sim$72B
10 perplexity/sonar-pro 90.32 196/217 70B
10 qwen/qwen3-next-80b-a3b-instruct 90.32 196/217 80B/3B
13 deepseek/deepseek-v3.1-terminus 89.86 195/217 671B/37B
13 qwen/qwen-plus 89.86 195/217 U
13 qwen/qwen3-next-80b-a3b-thinking 89.86 195/217 80B/3B
16 deepseek/deepseek-chat 89.40 194/217 671B/37B
16 o1 89.40 194/217 $\sim$200B
16 qwen/qwen-max 89.40 194/217 U
19 anthropic/claude-3-opus 88.94 193/217 U
20 openai/o1-mini 88.02 191/217 $\sim$100B
21 meta-llama/llama-3.1-405b-instruct 87.56 190/217 405B
21 qwen/qwen-2.5-72b-instruct 87.56 190/217 72B
23 mistralai/mistral-large 86.18 187/217 123B
24 amazon/nova-pro-v1 86.11 186/216 $\sim$90B
25 gpt-4-turbo 85.25 185/217 1.8T/280B
26 meta-llama/llama-3-70b-instruct 84.79 184/217 70B
26 meta-llama/llama-3.1-70b-instruct 84.79 184/217 70B
28 dziner-qwen-2.5-coder-32b 83.87 182/217 $\sim$32B
29 openai/gpt-4-0314 83.41 181/217 1.8T/280B
29 gpt-4-turbo-preview 83.41 181/217 1.8T/280B
31 amazon/nova-lite-v1 82.03 178/217 $\sim$20B
32 dziner-qwen-2.5-7b 81.57 177/217 $\sim$7B
32 gpt-4 81.57 177/217 1.8T/280B
34 openai/gpt-4o-mini 81.11 176/217 $\sim$8B
34 openrouter/auto 81.11 176/217 N/A
36 qwen/qwen-2.5-7b-instruct 79.72 173/217 7B
37 google/gemma-2-27b-it 79.26 172/217 27B
38 anthropic/claude-3.5-haiku 77.57 166/214 U
39 mistralai/mixtral-8x22b-instruct 76.81 159/207 141B/39B
40 mistralai/mistral-7b-instruct 75.35 162/215 7B
41 google/gemma-2-9b-it 75.12 163/217 9B
42 amazon/nova-micro-v1 74.65 162/217 $\sim$11B
42 mistralai/mistral-small 74.65 162/217 24B
44 anthropic/claude-3-haiku 74.04 154/208 $\sim$20B
45 openai/gpt-3.5-turbo-16k 72.60 151/208 $\sim$20B
46 mistralai/mixtral-8x7b-instruct 72.33 149/206 47B/13B
47 meta-llama/llama-3-8b-instruct 71.89 156/217 8B
48 mistralai/pixtral-12b 71.83 153/213 12B
49 openai/gpt-3.5-turbo 70.67 147/208 $\sim$20B
50 meta-llama/llama-3.1-8b-instruct 69.12 150/217 8B
51 llava-v1.6-34b 66.82 145/217 34B
52 lmms-lab/llava-onevision-qwen2-7b-si 66.36 144/217 7B
53 lmms-lab/llava-onevision-qwen2-7b-ov-chat 65.90 143/217 7B
54 lmms-lab/llava-onevision-qwen2-7b-ov 65.44 142/217 7B
55 meta-llama/llama-3.2-3b-instruct 64.98 141/217 3B
56 arcee-ai/afm-4.5b 62.21 135/217 4.5B
57 mistralai/mistral-7b-instruct-v0.1 59.50 119/200 7B
58 perplexity/sonar 59.24 125/211 70B
59 undi95/remm-slerp-l2-13b 57.08 121/212 13B
60 llamat-3-chat 57.14 124/217 8B
61 alibaba/tongyi-deepresearch-30b-a3b 55.76 121/217 30B/3B
62 gryphe/mythomax-l2-13b 53.77 114/212 13B
63 llava-v1.6-mistral-7b 52.99 115/217 7B
64 llamat-2-chat 50.69 110/217 7B
65 lmms-lab/llava-onevision-qwen2-0.5b-ov 47.47 103/217 0.5B
66 llava-v1.5-13b 46.54 101/217 13B
67 lmms-lab/llava-onevision-qwen2-0.5b-si 46.08 100/217 0.5B
68 qwen/qwen3-coder-flash 43.46 83/191 U
69 qwen/qwen3-coder-plus 22.99 43/187 U
70 honeybee-13b 22.12 48/217 13B
71 qwen/qwen-2.5-coder-32b-instruct 21.03 45/214 32B
72 honeybee-7b 19.35 42/217 7B
73 llava-v1.5-7b 17.97 39/217 7B
74 llamat-2 16.59 36/217 7B
@@
Fig. 6: Comparison of closed-book performance by model size group. Large Models (>>70B parameters or advanced architectures like GPT-4, GPT-5, O1, O3-mini, n=27); Mid-sized Models (>>7B-70B parameters including LLaVA-34B, QWEN2-7B, Mistral-7B, various 8B-13B models, n=22); Small Models ($\leq$7B parameters including QWEN2-0.5B, LLaMA-3.2-3B, Honeybee-7B, LLaMAT-2, n=25). Analysis excludes unreliable llava-v1.6-vicuna variants.